I need to solve this problem: What is the next number in this sequence:
1,2,8,48,384. I have tried to identify a pattern and have not been able to.
It is multiple choice here are the choices:a. 786 b.1536 c. 700, d. 580, e. 3,840
I don't know how to solve this. I tried to look at division/multiplication, subtaction/addition to identify a pattern. Thank you.

Found 2 solutions by Cromlix, Alan3354:
Answer by Cromlix(4381)

Hi there,
1, 2, 8, 48, 384
2/1 = 2
8/2 = 4
48/8 = 6
384/48 = 8
So next number should be 3,840
3,840/384 = 10
Hope this helps:-)

Answer by Alan3354(69443)

What is the next number in this sequence:
1,2,8,48,384.
-------
2 = 1*2
8 = 2*4
48 = 8*6
384 = 48*8
-----------
The multipliers are 2, 4, 6, 8 ...
--> 384*10 = 3840